How many milliliters of a 8% alcohol solution must be added to 12 milliliters of a 33% alcholol solution to produce a 20% alcohol solution?

Answer by checkley77(12844)

.33*12+.08*X=.20(12+X)
3.96+.08X=2.4+.20X
.08X-.20X=2.4-3.96
-.12X=-1.56
X=-1.56/-.12
X=13 ML. OF 33% SOLUTION.
.33*12+.08*13=.20(12+13
3.96+1.04=.20*25
5=5